OpenStudy (anonymous):

The equation y = 16x + 560 describes the number of calories in a recipe, where y is the total number of calories and x is the number of teaspoons of sugar added. The table shows the number of calories in another recipe, where y is the total number of calories and x is the number of teaspoons of honey added. How many calories are in a teaspoon of sugar, and how many are in a teaspoon of honey? Teaspoons of honey 1 2 4 7 Total calories 504 568 696 888 A teaspoon of sugar has____calories. A teaspoon of honey has____calories.

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Produce a table: Teaspoons of honey 1 2 3 4 16(no.of teaspoons) 16 x 1 16 x2 16x3 16 x4 +560 calories +560 +560 ~+560 +560 ---------------------------------------------------------------------------

OpenStudy (anonymous):

i don't get that

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

plz help @Directrix

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

confused on this one

OpenStudy (chosenmatt):

@bohotness @satellite73 @CausticSyndicalist @lizz123 @katkipe @KamiBug

OpenStudy (bohotness):

yes

OpenStudy (bohotness):

add

OpenStudy (chosenmatt):

help her i don't get it

OpenStudy (anonymous):

what do you mean add

OpenStudy (anonymous):

btw if i don't answer right away its my computer acting stupid

OpenStudy (bohotness):

okay many=add

OpenStudy (anonymous):

so i have to add the calories

OpenStudy (chosenmatt):

yup

OpenStudy (chosenmatt):

i get it now..

OpenStudy (anonymous):

sorry to be a bother but how do i do that

OpenStudy (bohotness):

yw

OpenStudy (bohotness):

chosen you get it now?

OpenStudy (chosenmatt):

little bit not really

OpenStudy (anonymous):

how am i suppose to add the calories

OpenStudy (anonymous):

you mean like 1 + 504?

OpenStudy (anonymous):

sorry if i'm a bother

OpenStudy (chosenmatt):

your not stop saying that your never a bother XD

OpenStudy (chosenmatt):

lol

OpenStudy (anonymous):

XD sorry i always think i talk to much

OpenStudy (bohotness):

tes

OpenStudy (bohotness):

well kind of

OpenStudy (anonymous):

do you mind explaining how to add them?

OpenStudy (anonymous):

@Angel_Bear19 plz help me this is last one i need to do and really want to get it done

OpenStudy (anonymous):

someone please help i want to get this over with

OpenStudy (anonymous):

hey @sleepyjess can you help me i'm so confused

OpenStudy (sleepyjess):

haha I just helped someone with this exact same problem the other day

OpenStudy (anonymous):

seriously?!

OpenStudy (sleepyjess):

Yep!

OpenStudy (anonymous):

well can you help me plz

OpenStudy (sleepyjess):

yeah

OpenStudy (anonymous):

thanks gosh

OpenStudy (sleepyjess):

y=16c+560 is the number of calories for the sugar, so to find 1 teaspoon, substitute 1 for x

OpenStudy (sleepyjess):

*16x

OpenStudy (anonymous):

so it would be 16x1

OpenStudy (anonymous):

16x1=16 16+560=576

OpenStudy (sleepyjess):

yep

OpenStudy (sleepyjess):

then the honey you have a table

OpenStudy (anonymous):

well i did all of them like that and i got this.... y=16x2+560 y=592 y=16x4+560 y=624 y=16x7+560 y=672

OpenStudy (anonymous):

now i am stuck

OpenStudy (sleepyjess):

you only need it for 1 teaspoon.....

OpenStudy (anonymous):

ok so my teaspoon would be 575 right?

OpenStudy (anonymous):

A teaspoon of honey has 575 calories, right?

OpenStudy (anonymous):

but how do i get the sugar amount?

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

the toatal number of calories is given by the equation y = 16x + 560

OpenStudy (anonymous):

if x = no of teaspoons then: total no of calories in 1 teaspoon = 16 (1) + 560 = 16 + 560 = 576 total no. of calories in 2 teaspopons = 16(2) + 560 = 32 + 560 = 592 total no. of calories in 3 teaspoons = 16(3) + 560 = 48 + 560 = 608 total no. of calories in 4 teaspoons = 16(4) + 560 = 64 + 560 = 624